OpenStudy (anonymous):

Use quadratic formula to solve: (1/3)x^2+(1/21)x-1=0 simplify, use radicals as needed, express complex numbers in terms of i.

OpenStudy (saifoo.khan):

\[\huge \frac{1}{4}(-1-\sqrt{589}),\frac{1}{4}(-1+\sqrt{589})\]

OpenStudy (anonymous):

Multiply each term by 21 7x^2 +x -21 = 0 a =7, b=1, c = -21 b^2 -4ac = (1)^2-4(7)(-21) = 589 sqrt(b^2-4ac) = sqrt(589) x = (-1 +sqrt(598))/14 or ( -1 - sqrt(589))/14